# Fine Spectrum from Crude Analytic Bootstrap

### v1

@@ 1 Introduction
[…]
@@
This identity was applied to the $O(N)$ model in [21] and here we use a generalisation of it for operators other than the tilt. Operators carrying $O(N-1)$ charge $J$ can be viewed as part of a line bundle on the conformal manifold, specifically in the $J$’s tensor product of the tangent bundle, see [22, 23]. In terms of (1.3), replacing the operators at 0 and $\infty$ with charge $J$ operators, the identity still holds with a factor of $J$ on the right-hand side due to the number of allowed index contractions as in (5.1).
@@ 2 The magnetic line dCFT
[…]
where $a=1,\ldots,N-1$. The bulk action has $O(N)$ symmetry. This is broken to $O(N-1)$ by the choice of $\phi_{N}$ in the last term, which defines the line defect along $x^{\mu}(\tau)=(\tau,\mathbf{0})$. In the following we omit the index $N$ from $\phi_{N}$.
@@
Here labels $I,J$ are generalised indices that include the operators and their quantum numbers under the global symmetries $O(N-1)$ and $O(d-1)$. Then $\mathcal{T}_{IJ}$ identifies the primaries and has the appropriate group tensor structures. The scaling dimension of $\mathcal{O}_{I}$ is denoted by $\Delta_{I}$.
@@
is the conformally invariant cross-ratio. If we order the operators along the defect as $\tau_{1}<\tau_{2}<\tau_{3}<\tau_{4}$, then the cross-ratio satisfies $10<\chi<$ 1. The function $G_{IJKL}(\chi)$ can be expanded in s-channel conformal blocks as
@@
where $\Delta_{IJ}=\Delta_{I}-\Delta_{J}$ and similarly for $\Delta_{KL}$ and ${{}_{2}F_{1}}$ is a hypergeomentric function. For our purposes below we split the sum over $M$ in (2.9) into subspaces of operators ${\cal M}_{n,s,J}$ of fixed quantum numbers $s$ under $O(d-1)$, $J$ under $O(N-1)$ and order $\varepsilon^{0}$ dimension $n\in\mathbb{N}$. Thus,
@@ 3 Crude analytic bootstrap and unmixing
The four-point function of $\phi$ at linear order in $\varepsilon$ was calculated in [4]. The associated function $G(\chi)$ that features in (2.7) in this case is given by
@@
$\lambda_{n}^{2}$ =$=\begin{cases}\text{O}(\varepsilon^{2})\,,&\text{$n$ odd,}\\[2.0pt] \displaystyle\frac{n!(n-1)!}{(2n-3)!}\big{(}1-\big{(}4-4H_{n}+2\gamma_{n}(H_{2n-2}-H_{n-1})\big{)}\varepsilon\big{)}+\text{O}(\varepsilon^{2})\,,&\text{$n$ even,}\end{cases}$ (3.8)
@@
$\lambda_{n}^{2}$ =$=\begin{cases}\text{O}(\varepsilon^{2})\,,&\text{$n$ odd,}\\[2.0pt] \displaystyle\frac{n!(n-1)!}{(2n-3)!(N-1)}\big{(}1-2\gamma_{n}(H_{2n-2}-H_{n-1})\varepsilon\big{)}+\text{O}(\varepsilon^{2})\,,&\text{$n$ even.}\end{cases}$ (3.11)
@@
For this function, a procedure like the one described above is now unsuccessful if a single operator is assumed to be exchanged for each $n$. This indicates that there are in fact multiple physical operators that are exchanged, and that the results in (3.7), (3.8), (3.10) and (3.11) represent averages over the contributions of those operators. The number of physical operators with dimension close to $n$ depends on $n$, but we find that $G_{ab}(\chi)$ can be bootstrapped if we assume a minimum of two operators at each even $n$.
In fact, two operators at each even $n$ are enough to simultaneously bootstrap the singlet channels of all three correlators. For $n=2$, this bootstrap completely resolves the spectrum of singlets. So we turn now from examining single correlation functions to systems of multiple ones to focus on full sectors of exchanged operators in (2.11), starting with ${\mathcal{M}}_{2,\text{singlet},\text{singlet}}$.
@@
To understand the properties of these matrices and their submatrices, consider $M_{1}=\Lambda\Lambda^{T}$ a $k\times k$ matrix of rank $l$. It is symmetric and there exists an $l\times k$ matrix $\cal R$ such that ${\mathcal{R}}M_{1}{\mathcal{R}}^{T}=\mathds{1}_{l}$, the $l\times l$ identity matrix. One can also replace $\Lambda$ with a $k\times l$ matrix $L$ such that $M_{1}=LL^{T}$ and ${\mathcal{R}}L=\mathds{1}_{l}$. Then we can rewrite $M_{2}=\Lambda\Gamma\Lambda^{T}=LDL^{T}$ with $D={\mathcal{R}}M_{2}{\mathcal{R}}^{T}$ and clearly the eigenvalues of $\Gamma$ are the $l$ eigenvalues of $D$ and some extra eigenvalue that plays no role in $M_{2}$ and can be assumed to vanish. While this procedure works, in practice we actually do not need to find this particular ${\mathcal{R}}$. Instead we take any $l\times k$ matrix $R$ of rank $l$ such that the image of $R^{T}$ does not include the kernel of $\Lambda^{T}$. Under these conditions, $RM_{1}R^{T}=RL(RL)^{T}$ is an invertible $l\times l$ matrix and $(RM_{2}R^{T})(RM_{1}R^{T})^{-1}=(RL)D(RL)^{-1}$ is a full rank $l\times l$ matrix with the same eigenvalues as $RM_{2}R$, or the relevant eigenvalues of $\Gamma$. In the example above, $R$ is any of the $2\times 3$ matrices $R_{i}$ with $i=1,2,3$ defined such that their left action on a $3\times 3$ matrix removes the $i$-th row. In Section 4 we sometimes need to remove more than a single line.
@@
For the example presented above, we follow the latter approach and calculate one extra four-point function, e.g. $\langle\phi s_{-}\phi s_{-}\rangle$, which allows then to determine all three-point function coefficients to order $\varepsilon$. We computed this four-point function using Feynman diagrams and applying our analytic bootstrap procedure we find an exchanged operator of dimension close to one, which is simply $\phi$ in the ϕ×sitalic-$\phi\times s_{-}$ OPE. This gives us explicitly $\lambda_{\phi s_{-}\phi}^{2}=\lambda_{\phi\phi s_{-}}^{2}$ and with this information as well as the information contained in $\langle\phi\phi\phi\phi\rangle$, $\langle tttt\rangle$ and $\langle\phi\phi tt\rangle$, we obtain
@@ 4.1 Dimension $\Delta\sim 1$
At dimension close to one we have the operators made of the basic fields $\phi_{a}$ and $\phi$, both acquiring anomalous dimensions compared to the free-field value of $21-\varepsilon/$ 2 [1, 4]. The former are known as tilts and play an important role in the integral identities in Section 5 below. The details including their representation under $O(N-1)$ and the transverse rotation group $O(d-1)$ are in Table 1.
Name $O(N-1)$ $O(d-1)$ Dimension Fields
@@ 4.2 Dimension $\Delta\sim 2$
Name $O(N-1)$ $O(d-1)$ Dimension Fields
$s_{+}$, $s_{-}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $2+\frac{3N+20\pm\sqrt{N^{2}+40N+320}}{2(N+8)}\varepsilon$ $\phi\phi$, $\phi_{a}\phi_{a}$
@@ 4.3 Dimension $\Delta\sim 3$
Name $O(N-1)$ $O(d-1)$ Dimension Fields
@@
For the singlets at dimension $\Delta\sim 3$, which in the language of (2.11) is $\mathcal{M}_{3,S,S}$, the possible combinations of fields are $\phi\phi\phi$, $\phi_{a}\phi_{a}\phi$, $\partial_{\mu}\partial^{\mu}\phi$ and $\partial_{\tau}^{2}\phi$. The last one is a descendant of $\phi$ and the next to last is related to the first two operators by the equations of motion. To find the dimensions of the former two we note that they should appear in ϕ×s±italic-$\phi\times s_{\pm}$ and $t_{a}\times V_{a}$. The matrix of correlators which we consider (with repeated elements removed) is
@@
For the case $\mathcal{M}_{3,S,V}$ of the $O(N-1)$ vectors, there are four fundamental field combinations, up to total $\tau$-derivatives, namely $\phi\phi\phi_{a}$, $\phi_{b}\phi_{b}\phi_{a}$, $\partial_{\mu}\partial^{\mu}\phi_{a}$ and $\phi\partial_{\tau}\phi_{a}$. The fourth actually decouples from the other three since it transforms differently under parity and unlike the others appears in OPE channels where one would normally look for even dimensional $O(N-1)$ vector operators. In particular, this operator makes a contribution to $\phi\times t_{a}$, $V_{a}\times s_{\pm}$ and $V_{b}\times T_{ba}$. We computed the correlators
@@
Let us turn to sectors that have non-zero transverse spin under $O(d-1)$. There are two fundamental field combination that are $O(N-1)$ singlets and $O(d-1)$ vectors, namely $\phi\partial_{\mu}\phi$ and $\phi_{a}\partial_{\mu}\phi_{a}$. The corresponding operators belong to $\mathcal{M}_{3,V,S}$ and appear in the $O(N-1)$ singlet channels of $\phi\times D_{\mu}$ and $t_{a}\times U_{b\mu}$. The relevant correlators are
@@
For operators that transform as vectors under both $O(N-1)$ and $O(d-1)$, there are again two fundamental field realisations, $\phi\partial_{\mu}\phi_{a}$ and $\phi_{a}\partial_{\mu}\phi$. There is a linear combination of these that is a total derivative, but this is not a descendant since it is a transverse total derivative. The corresponding operators contribute to $\phi\times U_{a\mu}$ and $t_{a}\times D_{\mu}$, which means that the correlators to be considered are
@@
The remaining sectors are the antisymmetric and traceless-symmetric representations of $O(N-1)$ which also transform as vectors under $O(d-1)$. For both sectors, there is a single fundamental field combination, which stems from $\phi_{a}\partial_{\mu}\phi_{b}$ by either antisymmetrising the indices $a$ and $b$, or symmetrising and subtracting the trace. The corresponding operators appear in the antisymmetric and traceless-symmetric channels respectively of $t_{a}\times U_{b\mu}$. Therefore, these channels are the relevant components of the correlator $\langle t_{a}U_{b\mu}t_{c}U_{d\nu}\rangle$. We denote the antisymmetric operator by $A^{\prime}_{ab\mu}$ and the traceless-symmetric operator by $W_{ab\mu}$. Since both of these sectors have a single operator, there is no mixing and we can easily read off their OPE coefficients to $\text{O}(\varepsilon)$ as
@@
The remaining two sectors of operators with dimensions close to three both transform in the traceless-symmetric representation of $O(d-1)$, but transform as either a singlet $\mathcal{D}_{\mu\nu}$ or a vector $\mathcal{U}_{a\mu\nu}$ under $O(N-1)$. In terms of fundamental fields, these are given by $\partial_{\mu}\partial_{\nu}\phi-\operatorname{tr}$ and $\partial_{\mu}\partial_{\nu}\phi_{a}-\operatorname{tr}$ respectively. There is no OPE involving operators with dimensions close to one or two that contain contributions from these operators. They are therefore inaccessible to us through the use of our bootstrap methods. Instead, we calculated the dimensions via Feynman diagrams and found them to be $\Delta_{\mathcal{D}}=3-\frac{1}{5}\varepsilon$ and $\Delta_{\mathcal{U}}=3-\frac{2}{5}\varepsilon$.
@@ 4.4 Dimension $\Delta\sim 4$
Name $O(N-1)$ $O(d-1)$ Dimension Fields
@@
For the operators of dimension close to four, we unmix some specific sectors, including the singlet operators, of which there turns out to be eight. Possible fundamental field combinations are listed in the first two rows of Table 4. Among these field combinations, $\phi_{a}\phi_{a}\partial_{\tau}\phi$ is the only one with odd parity. The corresponding operator $p$ appears in ϕ×s±italic-$\phi\times s_{\pm}$ and $t_{a}\times V_{a}$, which means that the relevant correlators are
@@
The remaining even parity singlets appear in the singlet channels of the OPEs
$U_{a\mu}\times U_{b\nu}\,,\qquad$ $D_{\mu}\times D_{\nu}\,,$ (4.11)
$\phi\times\phi\,,\qquad$ $t_{a}\times t_{b}\,,$
$V_{a}\times V_{b}\,,\qquad$ $s_{+}\times s_{+}\,,\qquad$ $s_{+}\times s_{-}\,,\qquad$ $s_{-}\times s_{-}\,.$
@@
$D_{\mu}\times U_{a\nu}\,,\qquad$ $\phi\times t_{a}\,,$ (4.21)
$V_{a}\times s_{+}\,,\qquad$ $V_{a}\times s_{-}\,,\qquad$ $V_{b}\times T_{ba}\,.$
@@
The final two sectors that we unmix using our bootstrap methods are the parity odd and parity even two-index $O(N-1)$ traceless-symmetric operators. There are five such fundamental field expressions, one of which has odd parity, $\phi_{a}\phi_{b}\partial_{\tau}\phi-\operatorname{tr}$. The corresponding operator appears in $\phi\times T_{ab}$ and $t_{a}\times V_{b}$, which means that the relevant correlators are
@@
This system is degenerate, which indicates that there is a single operator $\varpi_{ab}$ involved. In addition to it’s dimension, which is written in Table 4, it is also possible to determine it’s OPE coefficients to $\text{O}(\varepsilon)$. We find them to be
@@
where $\langle U_{a\mu}U_{b\nu}U_{c\rho}U_{d\sigma}\rangle_{T;S}$ indicates the $O(N-1)$ traceless-symmetric $O(d-1)$ singlet channel of $\langle U_{a\mu}U_{b\nu}U_{c\rho}U_{d\sigma}\rangle$. This $5\times 5$ system is degenerate. If $U_{a\mu}\times U_{b\nu}$ is removed, then the resulting $4\times 4$ system is also degenerate. All other $4\times 4$ systems are non-degenerate, indicating four operators $\sigma^{4}_{M,ab}$ involved in the mixing problem. Their dimensions are

### v2

@@ 1 Introduction
${}R_{abcd}=2\int d\chi\log|\chi|$ $\langle t_{a}(1)t_{b}(\chi)t_{c}(0)t_{d}(\infty)\rangle_{c}=(\delta_{ac}\delta_{be}-\delta_{ae}\delta_{bc})\langle t_{e}(0)t_{d}(\infty)\rangle\,.$ (1.3)
@@
This identity was applied to the $O(N)$ model in [21] and here we use a generalisation of it for operators other than the tilt. Operators carrying O(N1)O(N-1) charge $J$ can be viewed as part of a line bundle on the conformal manifold, specifically in the $J$’s tensor product of the tangent bundle, see [22, 23]. In terms of (1.3), replacing the operators at 0 and $\infty$ with charge $J$ operators, the identity still holds with a factor of $J$ on the right-hand side due to the number of allowed index contractions as in (5.1).
@@ 2 The magnetic line dCFT
$S=\int d^{\hskip 0.5pt4-\varepsilon}x\,\big{(}\tfrac{1}{2}\partial^{\mu}\phi_{a}\partial_{\mu}\phi_{a}+\tfrac{1}{2}(\partial_{\mu}\phi_{N})^{2}+\tfrac{1}{4!}\lambda_{0}(\phi_{a}^{2}+\phi_{N}^{2})^{2}\big{)}+h_{0}\int_{-\infty}^{\infty}d\tau\,\phi_{N}(\tau,\mathbf{0})\,$, (2.1)
where $a=1,\ldots,N-1$. The bulk action has $O(N)$ symmetry. This is broken to O(N1)O(N-1) by the choice of $\phi_{N}$ in the last term, which defines the line defect along $x^{\mu}(\tau)=(\tau,\mathbf{0})$. In the following we omit the index $N$ from $\phi_{N}$.
@@
Here labels $I,J$ are generalised indices that include the operators and their quantum numbers under the global symmetries O(N1)O(N-1) and O(d1)O(d-1). Then $\mathcal{T}_{IJ}$ identifies the primaries and has the appropriate group tensor structures. The scaling dimension of $\mathcal{O}_{I}$ is denoted by $\Delta_{I}$.
@@
is the conformally invariant cross-ratio. If we order the operators along the defect as $\tau_{1}<\tau_{2}<\tau_{3}<\tau_{4}$, then the cross-ratio satisfies $0<\chi<1$. The function $G_{IJKL}(\chi)$ can be expanded in s-channel conformal blocks as
@@
where $\Delta_{IJ}=\Delta_{I}-\Delta_{J}$ and similarly for $\Delta_{KL}$ and $_{2}F_{1}$} is a hypergeomentric function. For our purposes below we split the sum over $M$ in (2.9) into subspaces of operators ${\cal M}_{n,s,J}$ of fixed quantum numbers $s$ under O(d1)O(d-1), $J$ under O(N1)O(N-1) and order $\varepsilon^{0}$ dimension $n\in\mathbb{N}$. Thus,
@@ 3 Crude analytic bootstrap and unmixing
The four-point function of $\phi$ (the singlet defect operator coming from the field $\phi_{N}$) at linear order in $\varepsilon$ was calculated in [4]. The associated function $G(\chi)$ that features in (2.7) in this case is given by
@@
$\lambda_{n}^{2}$ $=\begin{cases}\text{O}(\varepsilon^{2})\,,&\text{$n$ odd,}\\[2.0pt] \displaystyle\frac{n!(n-1)!}{(2n-3)!}\big{(}1-\big{(}4-4H_{n}+2\gamma_{n}(H_{2n-2}-H_{n-1})\big{)}\varepsilon\big{)}+\text{O}(\varepsilon^{2})\,,&\text{$n$ even,}\end{cases}$ (3.8)
@@
$\lambda_{n}^{2}$ $=\begin{cases}\text{O}(\varepsilon^{2})\,,&\text{$n$ odd,}\\[2.0pt] \displaystyle\frac{n!(n-1)!}{(2n-3)!(N-1)}\big{(}1-2\gamma_{n}(H_{2n-2}-H_{n-1})\varepsilon\big{)}+\text{O}(\varepsilon^{2})\,,&\text{$n$ even.}\end{cases}$ (3.11)
@@
For this function, a procedure like the one described above is now unsuccessful if a single operator is assumed to be exchanged for each $n\geqslant 2$. Indeed, assuming the exchange of a single operator $s$ of dimension close to two leads to
$\lambda_{\phi\phi s}\lambda_{tts}=0\,,\qquad\gamma_{s}\lambda_{\phi\phi s}\lambda_{tts}=\frac{2}{N+8}\,.$ (3.13)
Clearly these equations are inconsistent and this indicates that there are in fact multiple physical operators that are exchanged. Assuming the exchange of multiple operators $s_{i}$ has the effect of replacing $\lambda_{\phi\phi s}\lambda_{tts}\rightarrow\sum_{i}\lambda_{\phi\phi s_{i}}\lambda_{tts_{i}}$ and likewise for $\gamma_{s}\lambda_{\phi\phi s}\lambda_{tts}$, where the index $i$ runs over the exchanged operators. This tells us that the results in (3.7), (3.8), (3.10) and (3.11) represent averages over the contributions of multiple operators. In principle, we can assume the exchange of any arbitrarily large number of operators, but equations of this form can be made to be consistent by assuming the exchange of just two.
This is not surprising, as the number of physical scalars with dimension close to two is two. For $n>2$, there are more scalars; still, assuming two operators at each even $n$ is enough to simultaneously bootstrap the singlet channels of all three correlators. For $n=2$, this bootstrap completely resolves the spectrum of singlets. So we turn now from examining single correlation functions to systems of multiple ones to focus on full sectors of exchanged operators in (2.11), starting with ${\mathcal{M}}_{2,\text{singlet},\text{singlet}}$.
@@
[…]
@@
For the example presented above, we follow the latter approach and calculate one extra four-point function, e.g. $\langle\phi s_{-}\phi s_{-}\rangle$, which allows then to determine all three-point function coefficients to order $\varepsilon$. We computed this four-point function using Feynman diagrams and applying our analytic bootstrap procedure we find an exchanged operator of dimension close to one, which is simply $\phi$ in the $\phi\times s_{-}$ OPE. This gives us explicitly $\lambda_{\phi s_{-}\phi}^{2}=\lambda_{\phi\phi s_{-}}^{2}$ and with this information as well as the information contained in $\langle\phi\phi\phi\phi\rangle$, $\langle tttt\rangle$ and $\langle\phi\phi tt\rangle$, we obtain
@@ 4.1 Dimension $\Delta\sim 1$
At dimension close to one we have the operators made of the basic fields $\phi_{a}$ and $\phi$, both acquiring anomalous dimensions compared to the free-field value of $1-\varepsilon/2$ [1, 4]. The former are known as tilts and play an important role in the integral identities in Section 5 below. The details including their representation under O(N1)O(N-1) and the transverse rotation group O(d1)O(d-1) are in Table 1.
Name O(N1)O(N-1) O(d1)O(d-1) Dimension Fields
@@ 4.2 Dimension $\Delta\sim 2$
Name O(N1)O(N-1) O(d1)O(d-1) Dimension Fields
$_{+}, ss_{-}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $2+\frac{3N+20\pm\sqrt{N^{2}+40N+320}}{2(N+8)}\varepsilon$ $\phi\phi$, $\phi_{a}\phi_{a}$
@@ 4.3 Dimension $\Delta\sim 3$
Name O(N1)O(N-1) O(d1)O(d-1) Dimension Fields
@@
For the singlets at dimension $\Delta\sim 3$, which in the language of (2.11) is $\mathcal{M}_{3,S,S}$, the possible combinations of fields are $\phi\phi\phi$, $\phi_{a}\phi_{a}\phi$, $\partial_{\mu}\partial^{\mu}\phi$ and $\partial_{\tau}^{2}\phi$. The last one is a descendant of $\phi$ and the next to last is related to the first two operators by the equations of motion. To find the dimensions of the former two we note that they should appear in $\phi\times s_{\pm}$ and $t_{a}\times V_{a}$. The matrix of correlators which we consider (with repeated elements removed) is
@@
For the case $\mathcal{M}_{3,S,V}$ of the O(N1)O(N-1) vectors, there are four fundamental field combinations, up to total $\tau$-derivatives, namely $\phi\phi\phi_{a}$, $\phi_{b}\phi_{b}\phi_{a}$, $\partial_{\mu}\partial^{\mu}\phi_{a}$ and $\phi\partial_{\tau}\phi_{a}$. The fourth actually decouples from the other three since it transforms differently under parity and unlike the others appears in OPE channels where one would normally look for even dimensional O(N1)O(N-1) vector operators. In particular, this operator makes a contribution to $\phi\times t_{a}$, $V_{a}\times s_{\pm}$ and $V_{b}\times T_{ba}$. We computed the correlators
@@
Let us turn to sectors that have non-zero transverse spin under O(d1)O(d-1). There are two fundamental field combination that are O(N1)O(N-1) singlets and O(d1)O(d-1) vectors, namely $\phi\partial_{\mu}\phi$ and $\phi_{a}\partial_{\mu}\phi_{a}$. The corresponding operators belong to $\mathcal{M}_{3,V,S}$ and appear in the O(N1)O(N-1) singlet channels of $\phi\times D_{\mu}$ and $t_{a}\times U_{b\mu}$. The relevant correlators are
@@
For operators that transform as vectors under both O(N1)O(N-1) and O(d1)O(d-1), there are again two fundamental field realisations, $\phi\partial_{\mu}\phi_{a}$ and $\phi_{a}\partial_{\mu}\phi$. There is a linear combination of these that is a total derivative, but this is not a descendant since it is a transverse total derivative. The corresponding operators contribute to $\phi\times U_{a\mu}$ and $t_{a}\times D_{\mu}$, which means that the correlators to be considered are
@@
The remaining sectors are the antisymmetric and traceless-symmetric representations of O(N1)O(N-1) which also transform as vectors under O(d1)O(d-1). For both sectors, there is a single fundamental field combination, which stems from $\phi_{a}\partial_{\mu}\phi_{b}$ by either antisymmetrising the indices $a$ and $b$, or symmetrising and subtracting the trace. The corresponding operators appear in the antisymmetric and traceless-symmetric channels respectively of $t_{a}\times U_{b\mu}$. Therefore, these channels are the relevant components of the correlator $\langle t_{a}U_{b\mu}t_{c}U_{d\nu}\rangle$. We denote the antisymmetric operator by $A^{\prime}_{ab\mu}$ and the traceless-symmetric operator by $W_{ab\mu}$. Since both of these sectors have a single operator, there is no mixing and we can easily read off their OPE coefficients to $\text{O}(\varepsilon)$ as
@@
The remaining two sectors of operators with dimensions close to three both transform in the traceless-symmetric representation of O(d1)O(d-1), but transform as either a singlet $\mathcal{D}_{\mu\nu}$ or a vector $\mathcal{U}_{a\mu\nu}$ under O(N1)O(N-1). In terms of fundamental fields, these are given by $\partial_{\mu}\partial_{\nu}\phi-\operatorname{tr}$ and $\partial_{\mu}\partial_{\nu}\phi_{a}-\operatorname{tr}$ respectively. There is no OPE involving operators with dimensions close to one or two that contain contributions from these operators. They are therefore inaccessible to us through the use of our bootstrap methods. Instead, we calculated the dimensions via Feynman diagrams and found them to be $\Delta_{\mathcal{D}}=3-\frac{1}{5}\varepsilon$ and $\Delta_{\mathcal{U}}=3-\frac{2}{5}\varepsilon$.
@@ 4.4 Dimension $\Delta\sim 4$
Name O(N1)O(N-1) O(d1)O(d-1) Dimension Fields
@@
For the operators of dimension close to four, we unmix some specific sectors, including the singlet operators, of which there turns out to be eight. Possible fundamental field combinations are listed in the first two rows of Table 4. Among these field combinations, $\phi_{a}\phi_{a}\partial_{\tau}\phi$ is the only one with odd parity. The corresponding operator $p$ appears in $\phi\times s_{\pm}$ and $t_{a}\times V_{a}$, which means that the relevant correlators are
@@
The remaining even parity singlets appear in the singlet channels of the OPEs [^5] Note that contracting O(d1)O(d-1) indices gives $\delta^{\mu}{\!}_{\mu}=d-1=3-\varepsilon$.
$U_{a\mu}\times U_{b\nu}\,$$\qquad$ $D_{\mu}\times D_{\nu}\,$, (4.11)
$\phi\times\phi\,$$\qquad$ $t_{a}\times t_{b}\,$,
$V_{a}\times V_{b}\,$$\qquad$ $s_{+}\times s_{+}\,$$\qquad$ $s_{+}\times s_{-}\,$$\qquad$ $s_{-}\times s_{-}\,.$
@@
$D_{\mu}\times U_{a\nu}\,$$\qquad$ $\phi\times t_{a}\,$, (4.21)
$V_{a}\times s_{+}\,$$\qquad$ $V_{a}\times s_{-}\,$$\qquad$ $V_{b}\times T_{ba}\,.$
@@
The final two sectors that we unmix using our bootstrap methods are the parity odd and parity even two-index O(N1)O(N-1) traceless-symmetric operators. There are five such fundamental field expressions, one of which has odd parity, $\phi_{a}\phi_{b}\partial_{\tau}\phi-\operatorname{tr}$. The corresponding operator appears in $\phi\times T_{ab}$ and $t_{a}\times V_{b}$, which means that the relevant correlators are
@@
This system is degenerate, which indicates that there is a single operator $\varpi_{ab}$ involved. In addition to its dimension, which is written in Table 4, it is also possible to determine its OPE coefficients to $\text{O}(\varepsilon)$. We find them to be
@@
where $\langle U_{a\mu}U_{b\nu}U_{c\rho}U_{d\sigma}\rangle_{T;S}$ indicates the O(N1)O(N-1) traceless-symmetric O(d1)O(d-1) singlet channel of $\langle U_{a\mu}U_{b\nu}U_{c\rho}U_{d\sigma}\rangle$. This $5\times 5$ system is degenerate. If $U_{a\mu}\times U_{b\nu}$ is removed, then the resulting $4\times 4$ system is also degenerate. All other $4\times 4$ systems are non-degenerate, indicating four operators $\sigma^{4}_{M,ab}$ involved in the mixing problem. Their dimensions are